# Closing the Gap Between SGP4 and High-Precision Propagation via Differentiable Programming

### 26 Feb 2024 v2

@@ 2.2 Differentiable Programming
Automatic differentiation is a technique used for computing and evaluating derivatives of functions. It can be applied to any function that can be represented as a sequence of elementary operations, whose derivatives are known [29]. The basic idea of AD is to build a graph of elementary operations with known derivatives, which is then used (either in forward or reverse mode) to precisely compute the derivative at a given point. Before automatic differentiation, derivatives could either be computed manually, with a time-consuming and error-prone procedure, through symbolic differentiation, or through numerical differentiation. This last alternative is not ideal due to the presence of truncation and round-off errors, which might substantially jeopardize the accuracy of the found derivatives, if not chosen carefully. On the other hand, symbolic differentiation would allow to access the derivative of the function for the entire domain, but it can result in the size of the expressions growing exponentially (something known as "expression swell"), as the number of variables and operations is increased, due to the product rule of derivative [30]. Automatic differentiation solves these issues, by returning the exact value (up to machine precision) of the derivatives of a computer program evaluated at a certain point, without incurring exponentially growing expressions [31]. The drawback, besides the fact that it is not applicable for non-differentiable functions, is that this is only the derivative evaluated at a given point, and one does not have access to the symbolic form of the derivative across the entire domain. The differentiability constraint implies that the computer program needs to comply with certain requirements and does not admit discontinuities, singularities, or non-differentiable features. In our case, we have implemented dSGP4 using the Python library Pytorch [32], which is one of the most popular open-source ML libraries that supports AD in use in deep learning research and applications. To ensure the fidelity of dSGP4, our approach encompassed several key verification steps. First of all, we validated the dSGP4 model’s outputs against the original SGP4 model. This involved the design and execution of continuous integration tests (that can be found in the open-source library) to confirm that all mirrored functionalities faithfully reproduced the behavior of the original implementation. Our verifications also extended to the examination of automatic differentiation outputs for the input variables. To ensure correctness, we compared (up to local truncation errors caused by the finite difference technique) that the results obtained from dSGP4 automatic differentiation functionality almost match the ones found via finite differencing of the original SGP4.
@@ 4.1 Methods
The ML-dSGP4 model represents a new hybrid differentiable propagator that retains the computational advantages of SGP4 while enabling training and improvement with accurate data. The model can be configured to return either SGP4 predictions or activate input and output corrections using trained neural networks, offering a flexible solution.
@@
This setup can be easily extended for multiple propagation points, for instance, by taking the mean squared error as a loss function, and by setting up a stochastic gradient descent algorithm to update the parameters. The advantage of ML-dSGP4, compared to ML-SGP4 where only the output of the SGP4 propagator is corrected (proposed and studied for example in [7], and does not require differentiability of the SPG4 itself). In the next section, we will delve into some numerical experiments to showcase the use of ML-dSGP4 and to highlight the comparison between ML-dSGP4 and the NN corrector.
@@ 4.2 Experiments
Table 2: Baseline, NN corrector, and ML-dSGP4 RMSE values for the normalized state, the position, and velocity after 25 epochs, for training, validation, and test datasets.

### 27 Feb 2024 v3

@@ 2.2 Differentiable Programming
Automatic differentiation is a technique used for computing and evaluating derivatives of functions. It can be applied to any function that can be represented as a sequence of elementary operations, whose derivatives are known [29]. The basic idea of AD is to build a graph of elementary operations with known derivatives, which is then used (either in forward or reverse mode) to precisely compute the derivative at a given point. Before automatic differentiation, derivatives could either be computed manually, with a time-consuming and error-prone procedure, through symbolic differentiation, or through numerical differentiation. This last alternative is not ideal due to the presence of truncation and round-off errors, which might substantially jeopardize the accuracy of the found derivatives, if not chosen carefully. On the other hand, symbolic differentiation would allow to access the derivative of the function for the entire domain, but it can result in the size of the expressions growing exponentially (something known as "expression swell"), as the number of variables and operations is increased, due to the product rule of derivative [30]. Automatic differentiation solves these issues, by returning the exact value (up to machine precision) of the derivatives of a computer program evaluated at a certain point, without incurring exponentially growing expressions [31]. The drawback, besides the fact that it is not applicable for non-differentiable functions, is that this is only the derivative evaluated at a given point, and one does not have access to the symbolic form of the derivative across the entire domain. The differentiability constraint implies that the computer program needs to comply with certain requirements and does not admit discontinuities, singularities, or non-differentiable features. In our case, we have implemented dSGP4 using the Python library Pytorch [32], which is one of the most popular open-source ML libraries that supports AD in use in deep learning research and applications. To ensure the fidelity of dSGP4, our approach encompassed several key verification steps. First of all, we validated the dSGP4 model’s outputs against the "Revisiting Spacetrack Report #3" SGP4 model [10]. This involved the design and execution of continuous integration tests (that can be found in the open-source library) to confirm that all mirrored functionalities faithfully reproduced the behavior of the original implementation [^4] https://github.com/esa/dSGP4/tree/master/tests, date of access: February 2024.. As a part of these tests, we also verified that we could reproduce the output of the Space Track official SGP4 release [^5] https://www.space-track.org/documentation#/sgp4, date of access: February 2024.: this was done by propagating several different TLEs at different future times (between a few hours to a few days), and checking that the results of the propagation match up to below 1 mm in position and 1 mm/s in velocity the official release results. Our verifications also extended to the examination of automatic differentiation outputs for the input variables. To ensure correctness, we compared (up to local truncation errors caused by the finite difference technique) that the results obtained from dSGP4 automatic differentiation functionality match the ones found via finite differencing of the original SGP4.
@@ 4.1 Methods
The ML-dSGP4 model represents a new hybrid differentiable propagator that retains the computational advantages of SGP4 while enabling training and improvement with accurate data. The model can be configured to return either SGP4 predictions or activate input and output corrections using trained neural networks.
@@
This setup can be easily extended for multiple propagation points, for instance, by taking the mean squared error as a loss function, and by setting up a stochastic gradient descent algorithm to update the parameters. The advantage of ML-dSGP4 lies in its capacity to learn both the initial neural network weights and SGP4 parameters, therefore offering a flexible solution. This differs from ML-SGP4, where only the output of the SGP4 propagator is corrected, as explored in prior research [7]. In the next section, we will delve into some numerical experiments to showcase the use of ML-dSGP4 and to highlight the comparison between ML-dSGP4 and the ML-SGP4.
@@ 4.2 Experiments
Table 2: Baseline, ML-SGP4, and ML-dSGP4 RMSE values for the normalized state, the position, and velocity after 25 epochs, for training, validation, and test datasets.